# MAT-DiSMech: A Discrete Differential Geometry-based Computational Tool for Simulation of Rods, Shells, and Soft Robots

### v1

@@ 8.  Conclusions
The MAT-DiSMech simulation framework presented in this paper offers a comprehensive and versatile approach to simulating soft robotic systems, specifically those composed of rods, shells, or their combinations. By leveraging Discrete Differential Geometry (DDG) methods, our MATLAB-based simulator achieves a balance between accuracy and efficiency, effectively addressing the challenges of simulating nonlinear dynamics inherent in soft robotics. The modular framework supports a broad range of soft robotic configurations and dynamic forces, including gravity, contact, Coulomb and viscous friction, and aerodynamic drag—allowing researchers to adapt the simulation to specific system requirements. We provide numerous showcase simulations for rod and shell-like structures, including an example involving contact with the ground, and demonstrate numerical validation against theoretical solutions. Furthermore, since controlled actuation of the soft body is important in soft robotics, the ability to simulate actuation through time-dependent adjustments of natural curvature and twist, highlights MAT-DiSMech’s adaptability for dynamic applications such as shape morphing, as is evident from the showcase simulation of a manta ray. In conclusion, MAT-DiSMech not only enhances the potential for rapid prototyping and validation in soft robotics but also bridges the gap between simulation and real-world performance, advancing research in Sim2Real pathways. In the future, we plan on expanding the scope of MAT-DiSMech by integrating data-driven modeling techniques to model material nonlinearity and black-box external forces.
@@ Appendix A Mid-edge Normal Bending
[…]
@@
[…]
@@
Appendix B Aerodynamic Drag Jacobian Expressions
@@
$\frac{\partial\mathbf{n}_{t}}{\partial\mathbf{x}_{i}}=\frac{1}{||\mathbf{n}_{t}||^{3}}\left[||\mathbf{n}_{t}||^{2}\mathbb{I}-\mathbf{n}_{t}\mathbf{n}_{t}^{\top}\right][(\mathbf{x}_{k}-\mathbf{x}_{j})\times]$
where $\mathbb{I}$ is the square identity matrix of size $(3\times 3)$ and $[\mathbf{p}\times]$ denotes the $(3\times 3)$ skew-symmetric cross product matrix for a vector $\mathbf{p}$. $\frac{\partial\mathbf{n}_{t}}{\partial\mathbf{x}_{n}}=0$ if $n\not\in\{i,j,k\}$.
The expression for $\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}$ stays the same as in the case of the viscous force Jacobian calculation.
$\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}=\begin{cases}\frac{\mathbb{I}}{dt}&i=j,\\ 0&i\neq j\end{cases}$
where $\mathbb{I}$ is the square identity matrix of size $(3\times 3)$.

### v2

@@ 7.2.  Validation for Cantilever Beam Case
To examine a limitation of the hinge-based bending model for shells, we perform simulations assessing its sensitivity to mesh quality. It is well known that the hinge model performs reliably only on meshes composed of nearly equilateral triangles with good aspect ratios [24], whereas ill-conditioned meshes with skewed elements or extreme aspect ratios can result in significant errors or convergence issues. Figure 12 presents the normalized deflection error for a cantilever beam under gravity, comparing the hinge and mid-edge bending models across five mesh configurations, from highly regular (equilateral) to severely distorted (non-uniform). The results indicate that the mid-edge bending model maintains consistently low errors across all mesh types, demonstrating robustness against mesh irregularities. In contrast, the hinge model shows significant errors for the right-isosceles and non-uniform meshes. The right-isosceles mesh features many hinges oriented obliquely to the main bending direction, while the non-uniform mesh contains triangles with poor aspect ratios and skewness. Overall, these results highlight the hinge method’s mesh dependence, whereas the mid-edge method’s edge-associated degrees of freedom substantially improve robustness to mesh quality.
Figure 12: Normalized deflection at the free end of a cantilever beam (length 0.1 m, width 0.02 m, thickness 0.001 m, $E=2$ GPa, $\nu=0.5$) under gravity, modeled with hinge and mid-edge bending across different mesh types: equilateral, random, right-isosceles, equilateral-aligned, and non-uniform.
@@ 8.  Conclusions
The MAT-DiSMech simulation framework presented in this paper offers a comprehensive and versatile approach to simulating soft robotic systems, specifically those composed of rods, shells, or their combinations. By leveraging Discrete Differential Geometry (DDG) methods, our MATLAB-based simulator achieves a balance between accuracy and efficiency, effectively addressing the challenges posed by geometrically nonlinear dynamics in soft robots. The modular design supports a broad range of soft robotic configurations and dynamic forces, including gravity, contact, Coulomb and viscous friction, as well as hydrodynamic and aerodynamic drag, allowing researchers to tailor simulations to specific system requirements. We present a diverse set of showcase examples involving complex environmental interactions, such as frictional ground contact and viscous fluid forces, for both rod- and shell-like structures. We also demonstrate numerical validation against theoretical solutions, highlighting the strengths and limitations of different modeling techniques.
Since controlled actuation plays a central role in soft robotics, MAT-DiSMech supports time-dependent adjustments of natural stretch, curvature, and twist. This enables the simulation of dynamic behaviors such as shape morphing, as demonstrated in examples including an earthworm, manta ray, snake, PneuNet actuator, and gripping manipulator. In conclusion, MAT-DiSMech provides an accurate and efficient platform for simulating soft robot dynamics. It facilitates rapid prototyping and validation of diverse soft robotic designs, thereby advancing research along Sim2Real pathways. In the future, we plan on expanding the scope of MAT-DiSMech by integrating data-driven modeling techniques to model material nonlinearity and black-box external forces.
@@ Appendix A Mid-edge Normal Bending
$\frac{\partial^{2}(\mathrm{Tr}(\Lambda^{2}))}{\partial\xi^{i}\partial\xi^{j}}$ $=2\hskip 1.0ptc^{i}\hskip 1.0ptc^{j}\hskip 1.0pts^{i}\hskip 1.0pts^{j}\hskip 1.0pt(\mathbf{t}^{i}\cdot\mathbf{t}^{j})^{2},$
@@
$\frac{\partial^{2}(\mathrm{Tr}(\Lambda))^{2}}{\partial\xi^{i}\partial\xi^{j}}$ $=2\hskip 1.0ptc^{i}\hskip 1.0ptc^{j}\hskip 1.0pts^{i}\hskip 1.0pts^{j}\hskip 1.0pt{l^{i}}^{2}{l^{j}}^{2},$
@@
Appendix B RFT Jacobian Expressions
For the Jacobian of the RFT hydrodynamic force, we provide the expressions for $\frac{\partial\hat{\mathbf{e}}^{k}}{\partial\mathbf{x}_{j}}$ and $\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}$ used in equation (30).
[…]
where $n$ can be any integer and $\mathbb{I}$ is the square identity matrix of size $(3\times 3)$.
$\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}=\begin{cases}\frac{\mathbb{I}}{dt}&i=j,\\ 0&i\neq j\end{cases}$ (34)
where $\mathbb{I}$ is the square identity matrix of size $(3\times 3)$.
Appendix C Aerodynamic Drag Jacobian Expressions
@@
[…]
where $\mathbb{I}$ is the square identity matrix of size $(3\times 3)$ and $\mathbf{p}\times]$ denotes the $(3\times 3)$ skew-symmetric cross product matrix for a vector $\mathbf{p}$. $\frac{\partial\mathbf{n}_{t}}{\partial\mathbf{x}_{n}}=0$ if $n\not\in\{i,j,k\}$.
The expression for $\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}$ stays the same as in (34).